OpenStudy (anonymous):

What is (3x)(12x^2) A. 15x^2 B. 15x^3 C. 36x^2 D. 36x^3

OpenStudy (anonymous):

multiply constant with constaant and x with x also use power law x^m.x^n=x^m+n

OpenStudy (anonymous):

so as (3times12)x^(1+2)=36x^3 see it//

OpenStudy (anonymous):

get it.?

OpenStudy (anonymous):

noit reallty

OpenStudy (anonymous):

what thing confusing..??

OpenStudy (anonymous):

wait is it d? i tried

OpenStudy (anonymous):

yes,its D..

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you are welcome..

OpenStudy (anonymous):

its not D...

OpenStudy (anonymous):

its D.right answer...

OpenStudy (anonymous):

how you say its not D..tell i will clear it to you..

OpenStudy (anonymous):

OH it is CX